OpenStudy (anonymous):

write the expression, using rational exponents, then simplify \[\sqrt[3]{125p^9q^6}\]

OpenStudy (mertsj):

Just make the top number the power and the bottom number the root.

OpenStudy (campbell_st):

this can be written as \[\sqrt[3]{(5)^3(p^3)^3(q^2)^3} = 5p^3q^2\]

OpenStudy (mertsj):

\[125^{\frac{1}{3}}\]

OpenStudy (mertsj):

Now you finish it.